OpenStudy (anonymous):

a plant grew 1/8 inch last week and 3/8 inch this week how many inches did the plant grow in the last two weeks

OpenStudy (anonymous):

options a. 1/2 b. 1/4 c. 5/8 d. 1/8

OpenStudy (sushi121212):

You have to add the 2 numerators together, but not the denominators. The denominators change only if it's a division or multiplication question: \[\frac{ 1 }{ 8 }+\frac{ 3 }{ 8 }=\frac{ 4 }{ 8 }\] Divide it by the common factor (which is 4) to get: \[\frac{ 1 }{ 4 }\]